OpenStudy (toxicsugar22):

Delbert prepares a 25% glugose solution of by mixing 2 ml of glucose with 8 ml of water. If he adds x ml of glucose to the solution, its concentration is given by C(X)=2+x/8+x toxicsugar22: How many ml of glucose should delberrt add to increase the concentration to 50%

OpenStudy (anonymous):

first find how much glucose it has by setting C(x)=.25\[.25 = {(2+x) \over (8+x)}\] find how much glucose does a 50% concentration solution has by setting the equation equal to 50%:\[.50 = {(2+x) \over (8+x)}\] Last, subtract the 50% x from the 25% x and you will have how much is need to add. is that right?

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

thamks

OpenStudy (anonymous):

np

OpenStudy (toxicsugar22):

can u help me with on more

OpenStudy (anonymous):

sure, put it up

OpenStudy (toxicsugar22):

The cost in thousands of dollers for extracting p% of a precious ore from a mine is given by the equation C(p)=360p/100-p Complete this showing the cost of extracting various percents of the ore . if 15 is P then what is C

OpenStudy (toxicsugar22):

do u get it

OpenStudy (anonymous):

don't yu just substitu 15 for P in the C(p) equation?

OpenStudy (toxicsugar22):

not in the 360p/100-p

OpenStudy (toxicsugar22):

ok so how do i do that

OpenStudy (anonymous):

what is 15 ? p=15% ?

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

so how do i do that can u show me please

OpenStudy (anonymous):

the we know that p =0.15 right?

OpenStudy (anonymous):

if we know that the cost is: C(p)=360p/100-p what would it be if p = 0.15? C(0.15) = ?

OpenStudy (toxicsugar22):

yes can u show me this one time

OpenStudy (anonymous):

C(0.15) = 360(0.15)/100-0.15

OpenStudy (toxicsugar22):

yes but it said do not convert percents to decimals

OpenStudy (toxicsugar22):

ok i got it

OpenStudy (anonymous):

sorry didn't see that 15/100 then

OpenStudy (toxicsugar22):

ok i got it

OpenStudy (toxicsugar22):

with that question what percent of the ore can be extracted if 540,000 can be spent on the extraction?

OpenStudy (anonymous):

now you set C(p) = 540,000 so, 540,000 = 360p/100-p solve for p

OpenStudy (toxicsugar22):

can u show me pleae

OpenStudy (anonymous):

540000(100-p) = 360p 54,000,000 - 540,000p = 360p 54,000,000 = 540,360p p = 99.93

OpenStudy (toxicsugar22):

thanks

OpenStudy (anonymous):

np

OpenStudy (toxicsugar22):

A computer store sells approximately 300 of its most popular model per year. The manager would like to minimize her annual inventory cost by ordering the optimal numbers of computers, x, at regular intervals. If she orders x computers in each shipment, the cost of storage willl be 6x dollers, and the cost of reording will be 300/x (15x+10) dollers. The inventory cost is the sum of the storage cost and the reording cost. use the distributive law to simplfy the expression for the rerding cost. Then express the inventory cost C as a function of x

OpenStudy (toxicsugar22):

can u help me